OpenStudy (anonymous):

Write 32x5y5 with only one exponent. Use parentheses.

OpenStudy (diyadiya):

2^5 =32 so \[\Huge 2^5 x^5 y^5 = (2xy)^5\]

OpenStudy (diyadiya):

Nope ,Why ?

OpenStudy (anonymous):

wouldn't it be 32(xy)^5 ?

OpenStudy (diyadiya):

32 (xy)^2 = 32x^2y^2

OpenStudy (anonymous):

sorry typo

OpenStudy (diyadiya):

lol ok

OpenStudy (anonymous):

32(xy)^5 would be right too though? cause it doesn't say simplify, right?

OpenStudy (anonymous):

or (32xy)^5 ? :)

OpenStudy (diyadiya):

With only one exponent & put in parantheses ,so it should (2xy)^5 (32xy)^5 = (32)^% x^5 y^5 :D

OpenStudy (diyadiya):

Oops i meant *(32)^5 x^5 y^5

OpenStudy (anonymous):

Ahh I get it, thanks :) do you know which exponent law it's from, that changes the 32 to a 2^5?

OpenStudy (anonymous):

the final answer is (2xy)^5?

OpenStudy (diyadiya):

I don't think there is a law for that when you prime factorise 32 you can see that 32 is 2 x 2 x 2 x 2 x 2 which makes it 2^5 @Ansleefbaby Yes :)

OpenStudy (anonymous):

and you only do that so it has the same exponent as the others :) thanks a lot diyadiya

OpenStudy (diyadiya):

Yes :) You're Welcome !! :)

OpenStudy (diyadiya):

and also because they can be combined togarther and put inside a parentheses.

OpenStudy (anonymous):

(2xy)^5 isnt correct.

OpenStudy (diyadiya):

Why not ? I'm SURE it is :)

OpenStudy (anonymous):

My test said incorrect.

OpenStudy (diyadiya):

But thats right!

OpenStudy (diyadiya):

You can ask any other users as well

OpenStudy (anonymous):

try 32(xy)^5 :P

OpenStudy (diyadiya):

Yeah that might be correct too^^

OpenStudy (anonymous):

lol just for peace of mind

OpenStudy (diyadiya):

Haha =D